Department of Physics Syracuse University Syracuse NY 13244-1130

June 15, 1995

Nazanin Azarnia Department of Mathematics Santa Fe Community Colleg Gainesville FL 32606-6200

Dear Sir,

Enclosed please find four copies of my paper entitled ”The Hypercube and Elementary Calculus”, which I am submitting for publication as a Classroom Capsule in The College Mathematics Journal.

Sincerely,

Roberto B. Salgado